# Improved Regret Bounds for Gaussian Process Upper Confidence Bound in Bayesian Optimization

### v1

@@ Abstract
This paper addresses the Bayesian optimization problem (also referred to as the Bayesian setting of the Gaussian process bandit), where the learner seeks to minimize the regret under a function drawn from a known Gaussian process (GP). Under a Matérn kernel with a certain degree of smoothness, we show that the Gaussian process upper confidence bound (GP-UCB) algorithm achieves $\widetilde{O}(\sqrt{T})$ cumulative regret with high probability. Furthermore, our analysis yields $O(\sqrt{T\ln^{4}T})$ regret under a squared exponential kernel. These results fill the gap between the existing regret upper bound for GP-UCB and the best-known bound provided by Scarlett [43]. The key idea in our proof is to capture the concentration behavior of the input sequence realized by GP-UCB, enabling a more refined analysis of the GP’s information gain.
@@ 1 Introduction
  • Furthermore, for a squared exponential kernel, we establish $O\left(\sqrt{T\ln^{4}T}\right)$ cumulative regret of GP-UCB. This improves the existing $O\left(\sqrt{T\ln^{d+2}T}\right)$ upper bound provided by Srinivas et al. [48] for $d>3$.
@@ 2 Preliminaries
where $C=\sqrt{\frac{2}{\ln(1+\sigma^{-2})}}$. From Eqs. (7) and (8), we conclude that the regret upper bound of GP-UCB is $O\left(\sqrt{\beta_{T}T\gamma_{T}(\mathcal{X})}\right)$ with probability at least $1-\delta$. By combining the explicit upper bound of $\gamma_{T}(\mathcal{X})$ [48, 54], we also obtain $O\left(\sqrt{T\ln^{d+2}T}\right)$ and $\widetilde{O}\left(T^{\frac{\nu+d}{2\nu+d}}\right)$ regret upper bounds for SE and Matérn kernels, respectively.
@@ 3 Improved Regret Bound for GP-UCB
Suppose Assumptions 1 and 2. Set $k=k_{\mathrm{SE}}$ or $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>2$. Furthermore, assume that $d,\nu,\ell$, $r$ , and $\sigma^{2}$ are fixed constants. Fix any $\delta_{\mathrm{GP}}\in(0,1)$, and set the confidence width parameter $\beta_{t}$ of GP-UCB as defined in Eq. (6) with any fixed $\delta\in(0,1-\delta_{\mathrm{GP}})$. Then, with probability at least $1-\delta_{\mathrm{GP}}-\delta$, the cumulative regret of GP-UCB (Algorithm 1) satisfies
$R_{T}=\begin{cases}\widetilde{O}\left(\sqrt{T}\right)&\mathrm{if}~{}k=k_{\mathrm{Mat\acute{e}rn}}~{}\mathrm{with}~{}2\nu+d\leq\nu^{2}~{}\mathrm{and}~{}4\nu^{2}+2\nu d\geq d^{2},\\ O\left(\sqrt{T\ln^{4}T}\right)&\mathrm{if}~{}k=k_{\mathrm{SE}}.\end{cases}$ (9)
@@ 3.2 Proof of Theorem 3
We set $\mathcal{T}(\varepsilon)=\{t\in[T]\mid f(\bm{x}^{\ast})-f(\bm{x}_{t})>\varepsilon\}$ and $\mathcal{T}^{c}(\varepsilon)=[T]\setminus\mathcal{T}(\varepsilon)$ in the above definition. A key observation is that, if we set sufficiently small $\varepsilon$ depending on the constants in Lemma 2, the inputs $(\bm{x}_{t})$ in $R_{T}^{(2)}(\varepsilon)$ (namely, inputs $(\bm{x}_{t})$ such that $f(\bm{x}^{\ast})-f(\bm{x}_{t})\leq\varepsilon$ holds) are on the locally linear or quadratic region around the maximizer $\bm{x}^{\ast}$ due to conditions 1, 3, and 4 in Lemma 2. The formal descriptions are provided in Lemma 20 in Appendix C. This fact is originally leveraged in Scarlett [43] to analyze the successive elimination-based algorithm. In the analysis of GP-UCB, it enables us to analyze the behavior of the sub-input sequence $\{\bm{x}_{t}\mid f(\bm{x}^{\ast})-f(\bm{x}_{t})\leq\varepsilon\}$ through the regularity constants $c_{\mathrm{lin}}$ and $c_{\mathrm{quad}}$. Below, we formally give the upper bound for $R_{T}^{(2)}(\varepsilon)$.
@@
$~{}~{}~{}+2\sqrt{C\beta_{T}T}\sum_{i=1}^{\overline{i}}\sqrt{\max\left\{\gamma_{(T/2^{i-1})}(\mathcal{B}_{2}(c_{\mathrm{lin}}^{-1}\eta_{i};\bm{x}^{\ast})),\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)\right\}},$
@@
We give the full proof in Appendix A.1. Since $\overline{i}=O(\ln T)$, the dominant term in the above lemma is given as:
@@
$R_{T}^{(2)}(\varepsilon)=\begin{cases}\widetilde{O}(\sqrt{T})~{}&\mathrm{if}~{}k=k_{\mathrm{Mat\acute{e}rn}}~{}\mathrm{with}~{}2\nu+d\leq\nu^{2},\\ O\left(\sqrt{T\ln^{4}T}\right)~{}&\mathrm{if}~{}k=k_{\mathrm{SE}}.\end{cases}$ (14)
The full proof is given in Appendix A.2. The remaining interest is the upper bound of $R_{T}^{(1)}(\varepsilon)$. The definition of $R_{T}^{(1)}(\varepsilon)$ is the same as the lenient regret [8], which is known to be smaller than the original regret $R_{T}$ in GP-UCB. Although Cai et al. [8] studies the frequentist setting, their proof strategy is also applicable to the Bayesian setting as described in Section 3.4 in [8]. The following lemma provides the formal statement about the upper bound of $R_{T}^{(1)}$.
@@
Fix any $\varepsilon>0$. Suppose $k=k_{\mathrm{SE}}$ or $k=k_{\mathrm{Mat\acute{e}rn}}$ with $4\nu^{2}+2\nu d\geq d^{2}$. When running GP-UCB, the following inequality holds under event $\mathcal{A}$:
$R_{T}^{(1)}(\varepsilon)$ $=O\left(\sqrt{\beta_{T}N(\varepsilon)\gamma_{N(\varepsilon)}(\mathcal{X})}\right)=\begin{cases}\widetilde{O}\left(\sqrt{T}\right)&\mathrm{if}~{}k=k_{\mathrm{Mat\acute{e}rn}},\\ O\left(\sqrt{(\ln T)^{d+3}(\ln\ln T)}\right)&\mathrm{if}~{}k=k_{\mathrm{SE}},\end{cases}$
where $N(\varepsilon)=\frac{16C\beta_{T}\gamma_{T}(\mathcal{X})}{\varepsilon^{2}}$ and $C=2/\ln(1+\sigma^{-2})$.
We provide the proof in Appendix A.3 for completeness. For both kernels, the above theorems shows that $R_{T}^{(1)}(\varepsilon)$ is dominated by the upper bound of $R_{T}^{(2)}(\varepsilon)$ under the conditions in Theorem 3. Finally, we obtain the desired results by aggregating the inequalities in Lemmas 5 and 6.
@@ A.1 Proof of Lemma 4
$\leq 2c_{\mathrm{sup}}\overline{T}+\frac{\pi^{2}}{3}\left(\log_{2}T+1\right)+2\sqrt{C\beta_{T}T}\sum_{i=1}^{\overline{i}}\sqrt{\gamma_{(T/2^{i-1})}(\mathcal{X}^{(i)})}.$ (35)
The last part of the proof is to specify the radius of the ball $\mathcal{B}_{2}(\cdot;\bm{x}^{\ast})$ such that $\mathcal{X}^{(i)}$ is included in it.
@@
$\begin{split}R_{T}^{(2)}(\varepsilon)&\leq 2c_{\mathrm{sup}}\overline{T}+\frac{\pi^{2}}{3}\left(\log_{2}T+1\right)\\ &~{}~{}~{}+2\sqrt{C\beta_{T}T}\sum_{i=1}^{\overline{i}}\sqrt{\max\left\{\gamma_{(T/2^{i-1})}(\mathcal{B}_{2}(\eta_{i}c_{\mathrm{lin}}^{-1};\bm{x}^{\ast})),\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)\right\}}.\end{split}$ (38)
@@ A.2 Proof of Lemma 5
$\frac{2c_{\mathrm{quad}}\ell^{2}}{\overline{\eta}_{i}ec_{d}}\geq T_{i}^{1/4}$ $\Leftrightarrow\frac{2c_{\mathrm{quad}}^{2}\ell^{2}}{ec_{d}\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}T_{i}^{-1/2}\ln^{\frac{d+1}{2}}T_{i}}\geq T_{i}^{1/4}$ (94)
$\Leftrightarrow T_{i}^{1/4}\geq\frac{ec_{d}\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}\ln^{\frac{d+1}{2}}T_{i}}{2c_{\mathrm{quad}}^{2}\ell^{2}}$ (95)
$\Leftarrow T_{i}^{1/4}\geq\frac{ec_{d}\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}\ln^{\frac{d+1}{2}}T}{2c_{\mathrm{quad}}^{2}\ell^{2}}$ (96)
$\Leftrightarrow T_{i}\geq\left(\frac{ec_{d}\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}\ln^{\frac{d+1}{2}}T}{2c_{\mathrm{quad}}^{2}\ell^{2}}\right)^{4}.$ (97)
Hence, if $\overline{T}\geq\left(\frac{ec_{d}\widetilde{C}_{\mathrm{SE}}^{2}\beta_{T}\ln^{d+1}T}{2c_{\mathrm{lin}}^{2}\ell^{2}}\right)^{2}$, we have
@@
$\overline{T}$ $\geq\max\left\{\left(\frac{ec_{d}}{2c_{\mathrm{quad}}^{2}\ell^{2}}\right)^{4},~{}\left(\frac{ec_{d}}{2c_{\mathrm{lin}}^{2}\ell^{2}}\right)^{2}\right\}\widetilde{C}_{\mathrm{SE}}^{4}\beta_{T}^{2}\ln^{2(d+1)}T.$ (101)
@@
Finally, by noting $\overline{T}=O(\ln^{2d+4}T)$ and $\overline{i}=O(\ln T)$, we obtain the following result from Lemma 4:
$R_{T}^{(2)}(\varepsilon)=O\left(\ln^{2d+4}T+\sqrt{T\ln^{4}T}\right).$ (103)
Since $d$ is a fixed constant, the above equation implies $R_{T}^{(2)}(\varepsilon)=O(\sqrt{T\ln^{4}T})$. ∎
@@ A.3 Proof of Lemma 6
$|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|\leq\frac{16C\beta_{\widetilde{t}}\gamma_{|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|}(\mathcal{X})}{\varepsilon^{2}}\leq\frac{16C\beta_{T}\gamma_{T}(\mathcal{X})}{\varepsilon^{2}}=N(\varepsilon)$ (111)
for any $\varepsilon>0$. Hence,
@@
$\leq 2c_{\mathrm{sup}}\sqrt{\frac{2}{\varepsilon}}+\frac{\pi^{2}}{6}+2\sqrt{C\beta_{T}N(\varepsilon)\gamma_{N(\varepsilon)}(\mathcal{X})}$ (114)
for any $\varepsilon>0$. In the above expressions, the last inequality follows from Lemma 21 and Eq. (111). Under fixed $\varepsilon>0$, the above inequality implies $R_{T}^{(\varepsilon)}=O\left(\sqrt{\beta_{T}N(\varepsilon)\gamma_{N(\varepsilon)}(\mathcal{X})}\right)$. Since $\gamma_{T}(\mathcal{X})=O(\ln^{d+1}T)$ and $\gamma_{T}(\mathcal{X})=\widetilde{O}(T^{\frac{d}{2\nu+d}})$ under $k=k_{\mathrm{SE}}$ and $k=k_{\mathrm{Mat\acute{e}rn}}$, respectively, we have
$R_{T}^{(1)}(\varepsilon)=\begin{cases}O\left(\sqrt{(\ln^{d+3}T)(\ln\ln T)}\right)~{}~{}&\mathrm{if}~{}~{}k=k_{\mathrm{SE}},\\ \widetilde{O}\left(\sqrt{T^{\frac{d}{2\nu+d}+\frac{d^{2}}{(2\nu+d)^{2}}}}\right)~{}~{}&\mathrm{if}~{}~{}k=k_{\mathrm{Mat\acute{e}rn}}.\end{cases}$ (115)
Finally, we have
$\frac{d}{2\nu+d}+\frac{d^{2}}{(2\nu+d)^{2}}\leq 1$ $\Leftrightarrow\frac{d(2\nu+d)+d^{2}}{(2\nu+d)^{2}}\leq 1$ (116)
$\Leftrightarrow 2d^{2}+2\nu d\leq 4\nu^{2}+4\nu d+d^{2}$ (117)
$\Leftrightarrow d^{2}\leq 4\nu^{2}+2\nu d.$ (118)
The above equations and Eq. (115) imply $R_{T}^{(1)}(\varepsilon)=\widetilde{O}(\sqrt{T})$ under $k=k_{\mathrm{Mat\acute{e}rn}}$ with $4\nu^{2}+2\nu d\geq d^{2}$. ∎

### v2

@@ Abstract
This paper addresses the Bayesian optimization problem (also referred to as the Bayesian setting of the Gaussian process bandit), where the learner seeks to minimize the regret under a function drawn from a known Gaussian process (GP). Under a Matérn kernel with a certain degree of smoothness, we show that the Gaussian process upper confidence bound (GP-UCB) algorithm achieves $\widetilde{O}(\sqrt{T})$ cumulative regret with high probability. Furthermore, our analysis yields $O(\sqrt{T\ln^{2}T})$ regret under a squared exponential kernel. These results fill the gap between the existing regret upper bound for GP-UCB and the best-known bound provided by Scarlett [43]. The key idea in our proof is to capture the concentration behavior of the input sequence realized by GP-UCB, enabling a more refined analysis of the GP’s information gain.
@@ 1 Introduction
  • Furthermore, for a squared exponential kernel, we establish $O\left(\sqrt{T\ln^{2}T}\right)$ cumulative regret of GP-UCB. This improves the existing $O\left(\sqrt{T\ln^{d+2}T}\right)$ upper bound provided by Srinivas et al. [48] for any $d\geq 1$.
@@ 2 Preliminaries
where $C=\frac{2}{\ln(1+\sigma^{-2})}$. From Eqs. (7) and (8), we conclude that the regret upper bound of GP-UCB is $O\left(\sqrt{\beta_{T}T\gamma_{T}(\mathcal{X})}\right)$ with probability at least $1-\delta$. By combining the explicit upper bound of $\gamma_{T}(\mathcal{X})$ [48, 54], we also obtain $O\left(\sqrt{T\ln^{d+2}T}\right)$ and $\widetilde{O}\left(T^{\frac{\nu+d}{2\nu+d}}\right)$ regret upper bounds for SE and Matérn kernels, respectively.
@@ 3 Improved Regret Bound for GP-UCB
Suppose Assumptions 1 and 2. Set $k=k_{\mathrm{SE}}$ or $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>2$. Furthermore, assume that $d,\nu,\ell$, $r$ , and $\sigma^{2}$ are fixed constants. Fix any $\delta_{\mathrm{GP}}\in(0,1)$, and set the confidence width parameter $\beta_{t}$ of GP-UCB as defined in Eq. (6) with any fixed $\delta\in(0,1-\delta_{\mathrm{GP}})$. Then, with probability at least $1-\delta_{\mathrm{GP}}-\delta$, the cumulative regret of GP-UCB (Algorithm 1) satisfies [^2] Correction for $k=k_{\mathrm{Mat\acute{e}rn}}$: In the earlier version of the paper, we missed and omitted some steps in the original proof of the lenient regret by Cai et al. [8]. Since this omission necessitates an additional assumption: $4\nu^{2}+2\nu d\geq d^{2}$ in the earlier version of the paper, it has been corrected in the present version. We are grateful to Jonathan Scarlett for pointing this out.  [^3] Correction for $k=k_{\mathrm{SE}}$: In the earlier version of the paper, we obtained $O(\sqrt{T\ln^{4}T})$ upper bound for $k=k_{\mathrm{SE}}$; however, we can reduce additional multiplicative $O(\ln T)$ term by trivial modification of the proof in Lemma 4, and we correct it in this version. We are grateful to Shion Takeno for pointing this out.
$R_{T}=\begin{cases}\widetilde{O}\left(\sqrt{T}\right)&\mathrm{if}~{}k=k_{\mathrm{Mat\acute{e}rn}}~{}\mathrm{with}~{}2\nu+d\leq\nu^{2},\\ O\left(\sqrt{T\ln^{2}T}\right)&\mathrm{if}~{}k=k_{\mathrm{SE}}.\end{cases}$ (9)
@@ 3.2 Proof of Theorem 3
We set $\mathcal{T}(\varepsilon)=\{t\in[T]\mid f(\bm{x}^{\ast})-f(\bm{x}_{t})>\varepsilon\}$ and $\mathcal{T}^{c}(\varepsilon)=[T]\setminus\mathcal{T}(\varepsilon)$ in the above definition. A key observation is that, if we set sufficiently small $\varepsilon$ depending on the constants in Lemma 2, the inputs $(\bm{x}_{t})$ in $R_{T}^{(2)}(\varepsilon)$ (namely, inputs $(\bm{x}_{t})$ such that $f(\bm{x}^{\ast})-f(\bm{x}_{t})\leq\varepsilon$ holds) are on the locally linear or quadratic region around the maximizer $\bm{x}^{\ast}$ due to conditions 1, 3, and 4 in Lemma 2. The formal descriptions are provided in Lemma 20 in Appendix C. This fact is originally leveraged in [43] to analyze the successive elimination-based algorithm. In the analysis of GP-UCB, it enables us to analyze the behavior of the sub-input sequence $\{\bm{x}_{t}\mid f(\bm{x}^{\ast})-f(\bm{x}_{t})\leq\varepsilon\}$ through the regularity constants $c_{\mathrm{lin}}$ and $c_{\mathrm{quad}}$. Below, we formally give the upper bound for $R_{T}^{(2)}(\varepsilon)$.
@@
$~{}~{}~{}+\frac{2\sqrt{2C\beta_{T}T}}{\sqrt{2}-1}\max_{i\in[\overline{i}]}\sqrt{\max\left\{\gamma_{(T/2^{i-1})}(\mathcal{B}_{2}(c_{\mathrm{lin}}^{-1}\eta_{i};\bm{x}^{\ast})),\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)\right\}},$
@@
We give the full proof in Appendix A.1. Here, the dominant term in the above lemma is given as:
@@
$R_{T}^{(2)}(\varepsilon)=\begin{cases}\widetilde{O}(\sqrt{T})~{}&\mathrm{if}~{}k=k_{\mathrm{Mat\acute{e}rn}}~{}\mathrm{with}~{}2\nu+d\leq\nu^{2},\\ O\left(\sqrt{T\ln^{2}T}\right)~{}&\mathrm{if}~{}k=k_{\mathrm{SE}}.\end{cases}$ (14)
The full proof is given in Appendix A.2. The remaining interest is the upper bound of $R_{T}^{(1)}(\varepsilon)$. The definition of $R_{T}^{(1)}(\varepsilon)$ is the same as the lenient regret [8], which is known to be smaller than the original regret $R_{T}$ in GP-UCB. Although Cai et al. [8] studies the frequentist setting, their proof strategy is also applicable to the Bayesian setting as described in Section 3.4 in [8]. The following lemma provides the formal statement about the upper bound of $R_{T}^{(1)}(\varepsilon)$.
@@
Fix any $\varepsilon>0$. Suppose $k=k_{\mathrm{SE}}$ or $k=k_{\mathrm{Mat\acute{e}rn}}$. Then, when running GP-UCB, $R_{T}^{(1)}(\varepsilon)=\widetilde{O}(1)$ holds under event $\mathcal{A}$.
We provide the proof in Appendix A.3 for completeness. For both kernels, $R_{T}^{(1)}(\varepsilon)$ is dominated by the upper bound of $R_{T}^{(2)}(\varepsilon)$. Finally, we obtain the desired results by aggregating the inequalities in Lemmas 5 and 6.
@@ A.1 Proof of Lemma 4
$\leq 2c_{\mathrm{sup}}\overline{T}+\frac{\pi^{2}}{3}\left(\log_{2}T+1\right)+2\sqrt{C\beta_{T}T}\sum_{i=1}^{\overline{i}}\frac{1}{2^{(i-1)/2}}\sqrt{\gamma_{(T/2^{i-1})}(\mathcal{X}^{(i)})}$ (35)
$\leq 2c_{\mathrm{sup}}\overline{T}+\frac{\pi^{2}}{3}\left(\log_{2}T+1\right)+\frac{2\sqrt{2C\beta_{T}T}}{\sqrt{2}-1}\max_{i\in[\overline{i}]}\sqrt{\gamma_{(T/2^{i-1})}(\mathcal{X}^{(i)})}.$ (36)
The last line follows from $\sum_{i=1}^{\overline{i}}\frac{1}{2^{(i-1)/2}}\leq\sum_{i=1}^{\infty}\frac{1}{2^{(i-1)/2}}=\frac{1}{1-1/\sqrt{2}}=\frac{\sqrt{2}}{\sqrt{2}-1}$. The last part of the proof is to specify the radius of the ball $\mathcal{B}_{2}(\cdot;\bm{x}^{\ast})$ such that $\mathcal{X}^{(i)}$ is included in it.
@@
$\begin{split}R_{T}^{(2)}(\varepsilon)&\leq 2c_{\mathrm{sup}}\overline{T}+\frac{\pi^{2}}{3}\left(\log_{2}T+1\right)\\ &~{}~{}~{}+\frac{2\sqrt{2C\beta_{T}T}}{\sqrt{2}-1}\max_{i\in[\overline{i}]}\sqrt{\max\left\{\gamma_{(T/2^{i-1})}(\mathcal{B}_{2}(\eta_{i}c_{\mathrm{lin}}^{-1};\bm{x}^{\ast})),\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)\right\}}.\end{split}$ (39)
@@ A.2 Proof of Lemma 5
$\frac{2c_{\mathrm{quad}}\ell^{2}}{\overline{\eta}_{i}ec_{d}}\geq T_{i}^{1/4}$ $\Leftrightarrow\frac{2c_{\mathrm{quad}}\ell^{2}}{ec_{d}\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}T_{i}^{-1/2}\ln^{\frac{d+1}{2}}T_{i}}\geq T_{i}^{1/4}$ (95)
$\Leftrightarrow T_{i}^{1/4}\geq\frac{ec_{d}\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}\ln^{\frac{d+1}{2}}T_{i}}{2c_{\mathrm{quad}}\ell^{2}}$ (96)
$\Leftarrow T_{i}^{1/4}\geq\frac{ec_{d}\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}\ln^{\frac{d+1}{2}}T}{2c_{\mathrm{quad}}\ell^{2}}$ (97)
$\Leftrightarrow T_{i}\geq\left(\frac{ec_{d}\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}\ln^{\frac{d+1}{2}}T}{2c_{\mathrm{quad}}\ell^{2}}\right)^{4}.$ (98)
Hence, if $\overline{T}\geq\left(\frac{ec_{d}\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}\ln^{\frac{d+1}{2}}T}{2c_{\mathrm{quad}}\ell^{2}}\right)^{4}$, we have
@@
$\overline{T}$ $\geq\max\left\{\left(\frac{ec_{d}}{2c_{\mathrm{quad}}\ell^{2}}\right)^{4},~{}\left(\frac{ec_{d}}{2c_{\mathrm{lin}}^{2}\ell^{2}}\right)^{2}\right\}\widetilde{C}_{\mathrm{SE}}^{4}\beta_{T}^{2}\ln^{2(d+1)}T.$ (102)
@@
Finally, by noting $\overline{T}=O(\ln^{2d+4}T)$, we obtain the following result from Lemma 4:
$R_{T}^{(2)}(\varepsilon)=O\left(\ln^{2d+4}T+\sqrt{T\ln^{2}T}\right).$ (104)
Since $d$ is a fixed constant, the above equation implies $R_{T}^{(2)}(\varepsilon)=O(\sqrt{T\ln^{2}T})$. ∎
@@ A.3 Proof of Lemma 6
$|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|\leq\frac{16C\beta_{\widetilde{t}}\gamma_{|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|}(\mathcal{X})}{\varepsilon^{2}}\leq\frac{16C\beta_{T}\gamma_{|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|}(\mathcal{X})}{\varepsilon^{2}}$ (112)
for any $\varepsilon>0$. Furthermore,
@@
$\leq 2c_{\mathrm{sup}}\sqrt{\frac{2}{\varepsilon}}+\frac{\pi^{2}}{6}+2\sqrt{C\beta_{T}|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|\gamma_{|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|}(\mathcal{X})}$ (115)
for any $\varepsilon>0$. In the above expressions, the last inequality follows from Lemma 21. The remaining part of the proof is to substitute the quantity $|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|$ in Eq. (115) into its upper bound, which is deduced from Eq. (112) depending on the kernel.
For $k=k_{\mathrm{SE}}$.
Under $k=k_{\mathrm{SE}}$, we crudely take the upper bound of $|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|$ as
$|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|\leq\frac{16C\beta_{T}\gamma_{|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|}(\mathcal{X})}{\varepsilon^{2}}\leq\frac{16C\beta_{T}\gamma_{T}(\mathcal{X})}{\varepsilon^{2}}.$ (116)
The above upper bound implies $|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|=O\left(\beta_{T}\gamma_{T}(\mathcal{X})\right)$. Since $\gamma_{T}(\mathcal{X})=O(\ln^{d+1}T)$ under $k=k_{\mathrm{SE}}$, Eq. (115) implies
$R_{T}^{(1)}(\varepsilon)$ $\leq 2c_{\mathrm{sup}}\sqrt{\frac{2}{\varepsilon}}+\frac{\pi^{2}}{6}+O\left(\sqrt{\beta_{T}(\beta_{T}\gamma_{T}(\mathcal{X}))\ln^{d+1}(\beta_{T}\gamma_{T}(\mathcal{X}))}\right)$ (117)
$=O\left(\beta_{T}\sqrt{(\ln^{d+1}T)\ln^{d+1}(\ln^{d+2}T)}\right)$ (118)
$=O\left(\sqrt{(\ln T)^{d+3}(\ln\ln T)^{d+1}}\right)$ (119)
$=\widetilde{O}(1).$ (120)
For $k=k_{\mathrm{Mat\acute{e}rn}}$.
Set $C_{\mathrm{Mat}}>0$ as the constant such that the following inequality holds:
$\forall t\geq 2,\gamma_{t}(\mathcal{X})\leq C_{\mathrm{Mat}}t^{\frac{d}{2\nu+d}}\ln^{\frac{4\nu+d}{2\nu+d}}t.$ (121)
The existence of $C_{\mathrm{Mat}}$ is guaranteed by the upper bound of MIG established in Corollary 8. Then, if $|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|\geq 2$ holds, Eq. (112) implies
$|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|\leq\frac{16C\beta_{T}C_{\mathrm{Mat}}|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|^{\frac{d}{2\nu+d}}\ln^{\frac{4\nu+d}{2\nu+d}}|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|}{\varepsilon^{2}}$ (122)
$\Rightarrow|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|\leq\frac{16C\beta_{T}C_{\mathrm{Mat}}|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|^{\frac{d}{2\nu+d}}\ln^{\frac{4\nu+d}{2\nu+d}}T}{\varepsilon^{2}}$ (123)
$\Leftrightarrow|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|^{\frac{2\nu}{2\nu+d}}\leq\frac{16C\beta_{T}C_{\mathrm{Mat}}\ln^{\frac{4\nu+d}{2\nu+d}}T}{\varepsilon^{2}}$ (124)
$\Leftrightarrow|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|\leq\left(\frac{16C\beta_{T}C_{\mathrm{Mat}}\ln^{\frac{4\nu+d}{2\nu+d}}T}{\varepsilon^{2}}\right)^{1+\frac{d}{2\nu}}.$ (125)
Therefore, we have $|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|=\widetilde{O}(1)$ under fixed $\varepsilon$, $d$, and $\nu$. Hence, from Eq. (115), we obtain $R_{T}^{(1)}(\varepsilon)=\widetilde{O}(1)$. ∎